# Quantum Simulation of Spin-Boson Models with Structured Bath

### v1

@@ Results
Figure 5: Simulations of vibrationally-assisted energy transfer (VAET). $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$ are fixed. (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$. Solid (dashed) curves depict theoretical predictions of the target (noise-aware) model. Circles represent experimental data and error bars are standard deviation over 100 repetitions. In (a), bath temperature is fixed to 0 K ($\bar{n}=0$) and $\nu$ is varied. In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated. (c) Time evolution of the donor-state population for noisy VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K. Solid (dashed) curves depict theoretical predictions of the spin-boson (dephased spin-oscillator) model. Circles denote experimental data. For $\kappa=0\>\text{cm}^{-1}$, error bars indicate the standard deviation over 100 repetitions. For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
@@
We first explore the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength $\kappa=30\>\text{cm}^{-1}$ and various values of $\nu$. When $\nu$ satisfies the resonance condition of VAET, i.e., $\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer, energy transfer occurs Gorman et al. (2018). Otherwise, energy remains localized at the donor site. Figure 5a shows the experimental data for simulating this model for the resonant ($\nu=52$ and $104\>\text{cm}^{-1}$) and non-resonant ($\nu=78\>\text{cm}^{-1}$) cases.
Furthermore, in Fig. 5b, we simulate the impact of initial temperature on the resonant ($\nu=104\>\text{cm}^{-1}$) VAET. We simulate baths of two temperatures 0 K and 300 K, which correspond to initial average phonon number $\bar{n}=0$ and $2.0$, respectively. For the latter case, SDK operations with random phase are applied to prepare the thermal state (when averaged over many trials). At first, the energy transfer is faster for the case of higher temperature, as more vibrational quanta assist the energy transfer. However, at later times, high temperature leads to a rapid decay of oscillations, as a thermal state of large $\bar{n}$ is a highly mixed state which quickly loses the coherent behavior. This pattern is accurately captured by our trapped-ion simulations.
@@
Noise (dephasing or damping) can be added to the VAET model. In Fig. 5c we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, showing good match with theoretical predictions. However, as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al. (2018); Tamascelli et al. (2018). This aligns with the aforementioned expectation that our method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime. To summarize, the short-time dynamics between the two models match as the 2-point correlation functions are the same; however, the long-time dynamics deviate due to mismatch in the higher-order correlation functions sup . Indeed, for nonzero $\epsilon$, the steady-state donor populations are different between the dephased spin-oscillator model and the spin-boson model, as clearly shown by the $\kappa=20\>\text{cm}^{-1}$ curves.
Discussion and outlook
An immediate future direction is to observe the effects of noisy modes on the signature behaviors of small-scale molecular models using trapped-ion simulations Kang et al. (2024). Examples are: conical intersections Whitlow et al. (2023); Valahu et al. (2023), polarized-light-induced electron transfer Sun et al. (2023); Zhang et al. (2024), and VAET in the donor-bridge-acceptor system Li et al. (2021). This may involve performing operations on a qudit Sun et al. (2023) or more than one qubits. For the latter case, entangling operations between the qubits Mølmer and Sørensen (1999) need to be added to the Trotterized instruction of operations.
Our method of using randomized operations to simulate spin-boson models is limited to weak-coupling regime. However, simulating a bath of dephased oscillators may be of independent interest, as this can be thought as a natural extension from Gaussian ($\hat{L}\propto\hat{b}$ or $\hat{b}^{\dagger}$) to non-Gaussian ($\hat{L}\propto\hat{b}^{\dagger}\hat{b}$) bath models Linowski et al. (2022); Barthel and Zhang (2022). In such case, strong-coupling regime may be readily explored. For example, an interesting question is whether the phase transition from coherent oscillations to localization for sub-Ohmic spin-boson models Nalbach and Thorwart (2013); Duan et al. (2017); Otterpohl et al. (2022) can be observed using dephased spin-oscillator models that match the spectral densities. As evidenced in Fig. 4, this may be simulated by our method applied to a few tens of motional modes, which can be readily utilized in a chain of 10 to 20 ions.
Simulations of spin-boson models in strong-coupling regime can be performed using the damped spin-oscillator model, as its spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density Lemmer et al. (2018); Tamascelli et al. (2018). Damped oscillators can be realized using sympathetic cooling of the trapped ions’ motional modes Lemmer et al. (2018); Schlawin et al. (2021). This requires using a multi-species ion chain, such that the optical pumping during sympathetic cooling does not destroy the qubit states that encode the spins. If accomplished, this implementation of damping can be combined with our dephasing method to simulate unexplored models of open quantum systems. We note that while the role of electronic states’ dephasing in molecular energy transfer has been heavily studied Grad et al. (1988); Goldsmith et al. (2006); Plenio and Huelga (2008); Hayes et al. (2010); Moreira et al. (2020), the role of vibrational modes’ dephasing is relatively less explored Yoneda et al. (2021).
@@
An ultimate goal is to accurately simulate the quantum dynamics of various models in physical Caldeira and Leggett (1983); Weiss et al. (1987), chemical Garg et al. (1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024). Classical simulation approaches continue to improve with advancements in numerical methods, yet they encounter fundamental challenges when dealing with large quantum systems. These challenges stem from the exponentially large Hilbert space and the quantum entanglement between the components. Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method Somoza et al. (2019), a state-of-the-art method for simulating open quantum systems. For simulating a one-dimensional chain of $N_{s}$ electronic states, each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix Somoza et al. (2019). A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics. In comparison, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales as only $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al. (2024). Also note that classical simulation methods typically struggle with complicated connectivity between the components, while for trapped ions all-to-all connectivity is granted both between qubits and between qubits and motional modes.
Realizing accurate trapped-ion simulations of models with many spins and noisy bosonic modes requires overcoming more challenges, such as efficiently characterizing the motional-mode parameters Kang et al. (2023b) and significantly lowering the hardware’s noise level. Nonetheless, our demonstration of accurate simulations using controlled dissipation represents a key step forward in addressing these challenges. In particular, the accuracy of our results is enabled by carefully designed calibration schemes and precise phase tracking sup ; Lobser et al. (2023). Also, our method enables using motional dephasing noise for simulating open quantum systems, which significantly relaxes the limitation on the experiment time (for simulating the time evolution) of our simulations imposed by hardware noise.
We anticipate this work inspires future works on improving the scale and accuracy of programmable quantum simulations of open quantum systems. Furthermore, our approach of finding models and parameter regimes that allow utilizing the dominant hardware noise may be of interest for various other analog quantum simulation platforms.
@@
During the preparation of our manuscript we became aware of Ref. So et al. (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode. This complements our experiment on simulating spin-boson models with structured bath using randomized operations.
@@ Acknowledgements.
The authors thank Yichao Yu for his valuable discussions about the experimental setup, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work. This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No. OMA-2120757 through a seed grant (experimental concepts and validation). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (data analysis and validation).
@@ Methods
Figure 6: Experimental setup and instruction of operations. (a) Cartoon depiction of the experimental setup illustrates the orientation of components. The magnetic field is perpendicular to the trap surface, while the individual beam is perpendicular to the trap surface and the global beam is parallel to it. The differential wave vector $\Delta\textbf{k}$ is oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes. The global beam comprises two tones driving the blue (BSB) and red (RSB) sideband transitions, respectively. (b) Simplified energy levels of the 171Yb+ ion and the laser pulses used for implementing the on-resonance spin-dependent kick (SDK) operation. (c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$. Here, $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit, $\phi_{s}$ ($\phi_{m}$) is the spin (motion) phase of the laser beams, $\tilde{\Omega}$ is the Rabi frequency of the sideband transition, and $\delta_{m}$ is the motion detuning. (d) Instruction of operations for simulating the dephased spin-oscillator model, which approximates the spin-boson model. ‘H’ represents the Hadamard gate on the qubit. ‘Kick (heat)’ represents the SDK operations with random $\phi_{m}$ that prepare the thermal state on average. ‘SQ’ (‘Kick (coh/dep)’) represents the single-qubit rotations (SDK operations) that simulate the first (second) term in Eq. (5). Coherent (dephased) spin-oscillator model is simulated by using a fixed (randomly drawn) $\delta_{m}$ value at each step.
@@
The simulations of spin-boson models are performed using a linear chain of 171Yb+ ions confined in a micro-fabricated surface trap Revelle (2020), as illustrated in Fig. 6a. The spin states are encoded to the 2S1/2-orbital hyperfine energy levels (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$. The bosonic bath modes are encoded as the collective radial motional modes of the ion chain. Details of our experimental setup can be found in Ref. Wang et al. (2020).
@@
Through the sequential application of Doppler cooling, EIT cooling, and sideband cooling techniques, the zigzag motional mode of the ion chain can be cooled to $\bar{n}=0.036(16)$. Following this, a $370$ nm laser is utilized to pump the ions to their ground state |$\ket{1}$. Transitions between the qubit states are performed using stimulated Raman transitions, driven by a pair of laser beams Wang et al. (2020). Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling the execution of Raman transitions at specified frequency detunings and phases. Specifically, the spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig. 6b. Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$ (see Fig. 6b), and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al. (2005). Subsequent to the final operation, the qubit-state population is accessed via the state-dependent detection technique.
@@
where $\hat{\sigma}_{\Delta}(t)\equiv\hat{\sigma}_{+}e^{i\Delta t}+\hat{\sigma}_{-}e^{-i\Delta t}$ and $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit. The evolution with respect to $\hat{H}_{I}(t)$ is discretized into $N$ time steps, each consisting of a single-qubit rotation and a SDK operation that simulate the first and second term, respectively. Specifically, for the SDK operation at the $j$-th time step, the spin (motion) phase of the laser beams Lee et al. (2005) is set as $\Delta t_{j}$ ($\nu t_{j}$), where $t_{j}\equiv(j-\frac{1}{2})T/N\tau$ and $\tau$ is the duration of the operation for each time step. This is equivalent to setting the symmetric (anti-symmetric) detuning from the sideband transitions, referred to as the spin (motion) detuning, as $\Delta T/N\tau$ ($\nu T/N\tau$) at all time steps. This method can be straightforwardly extended to the case where multiple modes are used; however, the laser phase needs to be tuned to track the phase of each motional mode correctly sup .
The Trotterization method here enables simulating the VAET model on a single qubit. This is different from Ref. Gorman et al. (2018), which simulates the same model using continuous evolution of two ions induced by global laser beams. Thus, Trotterization using individually addressed beams is more efficient in terms of the required number of qubits. Also, our approach can be straightforwardly extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al. (2024).
@@
Randomizing the control parameters is a key ingredient of our simulations of spin-boson models. By taking an average of many coherent evolutions, each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al. (2017). Heating (dephasing) is used to tune the temperature (spectral linewidth) of the bath.
First, infinite-temperature heating, described by the pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$, is induced by applying SDK operations with stochastically varying motion phase $\phi_{m}$. Specifically, we apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation. These stochastic operations are applied on the initial qubit-mode composite state |$\ket{+,n=0}$; note that $\ket{+}$ is an eigenstate of the spin operator with $\phi_{s}=0$, and thus the qubit is decoupled from these operations. When averaged over many trials of coherent evolutions, each executed with a distinct set of $N^{\prime}$ random $\phi_{m}$ values, this procedure prepares the thermal state with an average phonon number $\bar{n}=N^{\prime}\Gamma^{\prime}\tau^{\prime}$. This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states. See the Supplemental Material sup for derivations.
For the experimental data in Fig. 1a and Fig. 5b, the average phonon number $\bar{n}$ of the initial thermal state is tuned by varying the duration $\tau^{\prime}$ of SDK operations, where the number of steps $N^{\prime}$ is fixed to 5. The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplemental Material sup .
@@
In practice, the dephasing strength $\Gamma$ is varied by tuning the standard deviation of the random offsets to $\delta_{m}$ at each Trotterization step. This enables tuning the FWHM of each Lorentzian peak in the bath’s spectral density. The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplemental Material sup .

### v2

@@ Results
Figure 5: Simulations of vibrationally-assisted energy transfer (VAET). We set $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$. (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$. Solid (dashed) curves depict theoretical predictions of the target (noise-aware) model. Circles represent experimental data and error bars are standard deviation over 100 repetitions. In (a), bath temperature is fixed at 0 K ($\bar{n}=0$) and $\nu$ is varied. In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated. (c) Time evolution of the donor-state population for noisy VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K. Solid (dashed) curves depict theoretical predictions of the spin-boson (dephased spin-oscillator) model. Circles denote experimental data. For $\kappa=0\>\text{cm}^{-1}$, error bars indicate the standard deviation over 100 repetitions. For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
@@
Figure 5a shows the experimental data for simulating the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength ($\kappa=30\>\text{cm}^{-1}$). When $\nu$ satisfies the resonance condition of VAET ($\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer), energy transfer occurs Gorman et al. (2018). Otherwise, energy remains localized at the donor site. Figure 5b, shows the impact of initial temperature of the oscillators on the resonant VAET ($\nu=104\>\text{cm}^{-1}$). We simulate baths of two temperatures 0 K and 300 K corresponding to initial average phonon number $\bar{n}=0$ and $2.0$, respectively (see Methods). The energy transfer is faster for higher temperature bath in the beginning, as more vibrational quanta assist the energy transfer. However, high temperature leads to a rapid decay of oscillations over time as the coherence is lost.
@@
In Fig. 5c, we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, to study the impact of dissipative coupling. We note that as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al. (2018); Tamascelli et al. (2018). This is consistent with the expectation that the method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime. While the short-time dynamics between the two models match as the 2-point correlation functions are the same, the long-time dynamics deviate due to mismatch in the higher-order correlation functions for higher coupling strengths sup .
Discussions and outlook
@@
Simulations of large-scale spin-boson models with strong coupling in the presence of dissipation is crucial yet challenging problem in physical Caldeira and Leggett (1983); Weiss et al. (1987); Schindler et al. (2022), chemical Garg et al. (1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems. While classical simulation approaches continue to improve with advancements in numerical methods, they encounter fundamental challenges when dealing with large quantum systems that span exponentially large Hilbert space MacDonell et al. (2021); Kang et al. (2024). Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method, a state-of-the-art method for simulating open quantum systems Somoza et al. (2019). For simulating a one-dimensional chain of $N_{s}$ electronic states each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix. A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics. For these types of problems, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales only as $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al. (2024). The current state-of-the-art trapped ion systems offer several dozens of highly coherent spins and bosonic modes that are very densely coupled with each other, the complete quantum description of which is totally intractable with classical computers of today Cetina et al. (2022); Chen et al. (2023b). Developing various control methods over these quantum systems to study important physical, chemical, biological and material systems can open up a new computational framework in science.
The methods developed in this work utilizes fully programmable control methods to simulate dissipative processes. This approach can readily be applied to existing simulations of chemical processes to include coupling to environment Whitlow et al. (2023); Valahu et al. (2023); Sun et al. (2023); Zhang et al. (2024); Li et al. (2021). The dephased oscillator method can be used to realize non-Gaussian bath models Linowski et al. (2022); Barthel and Zhang (2022), relevant for studying physical processes such as phase transitions in strongly-coupled spin-boson models Duan et al. (2017); Otterpohl et al. (2022) and role of bosonic dephasing in molecular energy transfer Grad et al. (1988); Goldsmith et al. (2006); Plenio and Huelga (2008); Hayes et al. (2010); Moreira et al. (2020); Yoneda et al. (2021). High accuracy simulations will require development of efficient characterization methods Kang et al. (2023b) and novel control protocols in the trapped ion system sup ; Lobser et al. (2023) to implement various interaction terms in the target Hamiltonian. Combined with damped spin-boson models that can be implemented using sympathetic cooling Lemmer et al. (2018); Schlawin et al. (2021), the long-chain trapped ion systems can provide a versatile platform for simulating dynamics of complex open quantum systems.
@@
During the preparation of our manuscript we became aware of a complementary work So et al. (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode.
@@ Acknowledgements.
The authors thank Yichao Yu for his valuable discussions about the phase tracking method, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work. This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No. OMA-2120757 (experimental concepts and validation). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (data analysis and validation).
@@ Methods
Figure 6: Experimental setup and instruction of operations. (a) Schematic of the experimental setup illustrates the orientation of relevant fields. The magnetic field and the direction of the individual beam is perpendicular to the trap surface, while the global beam propagates is parallel to it. The differential wave vector $\Delta\textbf{k}$ is therefore oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes. The global beam comprises two tones driving the blue (BSB) and red (RSB) sideband transitions, respectively. (b) Simplified energy levels of the 171Yb+ ion and the laser pulses used for implementing the on-resonance spin-dependent kick (SDK) operation. (c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of the operator $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$. Here, $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit, $\phi_{s}$ ($\phi_{m}$) is the spin (motion) phase of the laser beams, $\tilde{\Omega}$ is the Rabi frequency of the sideband transition, and $\delta_{m}$ is the motion detuning. (d) Instruction of operations for simulating the dephased spin-oscillator model, which approximates the spin-boson model. ‘H’ represents the Hadamard gate on the qubit. ‘Kick (heat)’ represents the SDK operations with random $\phi_{m}$ that prepare the thermal state on average. ‘SQ’ (‘Kick (coh/dep)’) represents the single-qubit rotations (SDK operations) that simulate the first (second) term in Eq. (5). Coherent (dephased) spin-oscillator model is simulated by using a fixed (randomly drawn) $\delta_{m}$ value at each step.
@@
The simulations of spin-boson models are performed using a linear chain of 171Yb+ ions confined in a micro-fabricated surface trap Revelle (2020), as illustrated in Fig. 6a. The spin states are encoded in the hyperfine energy levels in the 2S1/2 manifold (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$. The bosonic bath modes are encoded as the collective radial motional modes of the ion chain. Details of our experimental setup can be found in Ref. Wang et al. (2020).
@@
Through the sequential application of Doppler cooling, EIT cooling, and sideband cooling techniques, the zigzag motional mode of the ion chain can be cooled to $\bar{n}=0.036(16)$. Following this, a $370$ nm laser is utilized to pump the ions to their ground state |$\ket{1}$. Transitions between the qubit states are performed using stimulated Raman transitions, driven by a pair of laser beams Wang et al. (2020). Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling single-qubit gates: we apply a bit flip ($X$) gate to initialize the ion to |$\ket{0}$ state. The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig. 6b. Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$, and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al. (2005). Subsequent to the final operation, the qubit-state population is accessed via the state-dependent detection technique.
@@
where $\hat{\sigma}_{\Delta}(t)\equiv\hat{\sigma}_{+}e^{i\Delta t}+\hat{\sigma}_{-}e^{-i\Delta t}$ and $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit. The evolution with respect to $\hat{H}_{I}(t)$ is discretized into $N$ time steps, each consisting of a single-qubit rotation and a SDK operation that simulate the first and second term, respectively. Specifically, for the SDK operation at the $j$-th time step, the spin (motion) phase of the laser beams Lee et al. (2005) is set as $\Delta t_{j}$ ($\nu t_{j}$), where $t_{j}\equiv(j-\frac{1}{2})T/N\tau$ and $\tau$ is the duration of the operation for each time step. This is equivalent to setting the symmetric (anti-symmetric) detuning from the sideband transitions, referred to as the spin (motion) detuning, as $\Delta T/N\tau$ ($\nu T/N\tau$) at all time steps. This method can be straightforwardly extended to the case where multiple modes are used; however, the laser phase needs to be tuned to track the phase of each motional mode correctly sup . This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al. (2024).
@@
Randomizing the control parameters is a key ingredient in our approach to simulating dissipation mechanisms in spin-boson models. By taking an average of many coherent evolutions each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al. (2017). Heating (dephasing) is used to tune the temperature (spectral linewidth) of the bath.
First, the bosonic modes at finite temperature is created by modeling heating process where the modes interact with an infinite-temperature external bath for a finite duration, described by aa pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$. We apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and motion phase $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation. These stochastic operations are applied on the initial qubit-mode composite state |$\ket{+,n=0}$; note that $\ket{+}$ is an eigenstate of the spin operator with $\phi_{s}=0$, and thus the qubit is decoupled from these operations. When averaged over many trials of coherent evolutions, each executed with a distinct set of $N^{\prime}$ random $\phi_{m}$ values, this procedure prepares the thermal state with an average phonon number $\bar{n}=N^{\prime}\Gamma^{\prime}\tau^{\prime}$. This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states (derivation provided in Supplementary Material sup ). For the experimental data in Fig. 1a and Fig. 5b, the average phonon number $\bar{n}$ of the initial thermal state is tuned by varying the duration $\tau^{\prime}$ of SDK operations, where the number of steps $N^{\prime}$ is fixed to 5. The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplementary Material sup .
@@
In practice, the dephasing strength $\Gamma$ is varied by tuning the standard deviation of the random offsets to $\delta_{m}$ at each Trotterization step. This enables tuning the FWHM of each Lorentzian peak in the bath’s spectral density. The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplementary Material sup .